# RNNs, CNNs and Transformers in Human Action Recognition: A Survey and A Hybrid Model

### v1

@@ 2.1 From Vanilla RNN to Attention-based Transformers
One-to-many represents a specific category of RNNs that is characterized by its ability to produce multiple outputs based on a single input provided to the model. This type of RNN is particularly useful in applications like Image Captioning, where a fixed input size results in a series of data outputs.
The Many-to-one RNNs merge a sequence of inputs into a single output through a series of hidden layers that learn relevant features. An illustrative instance of this RNN type is Sentiment Analysis, where the model analyzes a sequence of text inputs and produces a single output indicating the sentiment expressed in the text.
@@
The Many-to-many RNNs are employed to generate a sequence of output data from a sequence of input units. It can be categorized into two subcategories: Equal Size and Unequal Size. In the Equal Size subcategory, the input and output layers have the same size, see Many-to-many architecture in Figure 3. Several research efforts have emerged to tackle the limitation of fixed-size input-output sequences in machine translation tasks, as they fail to adequately represent real-world requirements. The Unequal Size subcategory can handle different sizes of inputs and outputs. A practical application of the the Unequal Size subcategory can be observed in Machine Translation. In this scenario, the model generates a sequence of translated text outputs based on a sequence of input sentences. Unequal Size subcategory employs an encoder-decoder architecture, where the encoder adopts the Many-to-One architecture, and the decoder adopts the One-to-Many architecture. One notable contribution in this area was made by Kalchbrenner and Blunsom [19], who pioneered the approach of mapping the entire input sentence to a vector. This work is related to the study conducted by Cho et al. [20], although the latter was specifically utilized to refine hypotheses generated by a phrase-based system [21]. In this architecture, the encoder component plays a crucial role in transforming the inputs into a singular vector, commonly referred to as the context. This context vector, typically with a length of 256, 512, or 1024, encapsulates all the pertinent information detected by the encoder from the input sentence, which serves as the translation target, see Figure 4(a). Subsequently, this vector is passed on to the decoder, which generates the corresponding output sequence. It’s important to note that both the encoder and decoder components in this architecture are RNNs.
2.1.1 Attention
The evolution of attention mechanisms in neural networks represents a significant advancement in the field of deep learning, particularly in tasks related to NLP and machine translation. Initially introduced by Graves et al. [22], the concept of attention mechanisms was designed to enhance the model’s ability to focus on specific parts of the input sequence when generating an output, mimicking the human ability to concentrate on particular aspects of a task. This foundational work laid the groundwork for subsequent developments in attention mechanisms, providing a mechanism for models to dynamically assign importance to different parts of the input data.
Building on Graves’ initial concept, Bahdanau et al. [23] introduced the additive attention mechanism, which was specifically designed to improve machine translation. This approach computes the attention weights through a feed-forward neural network, allowing the model to consider the entire input sequence and determine the relevance of each part when translating a segment. This additive form of attention significantly improved the performance of sequence-to-sequence models by enabling a more nuanced understanding and alignment between the input and output sequences [21]. Following this, Luong et al. [24] proposed the multiplicative attention mechanism, also known as dot-product attention, which simplifies the computation of attention weights by calculating the dot product between the query and all keys. This method not only streamlined the attention mechanism but also offered improvements in computational efficiency and performance in various NLP tasks, marking a pivotal moment in the evolution of attention mechanisms from their inception to more sophisticated and efficient variants.
@@ 2.1.1 Attention
Figure 4: Sequence-To-Sequence RNN with and without attention mechanism.
@@
2.1.2 Self-Attention
To this point, attention mechanisms in sequence-transformation models have primarily relied on complex RNNs, featuring an encoder and a decoder, the most successful models in language translation yet. However, in 2017, Vaswani et al. [5] introduced a simple network architecture known as the Transformer, see Figure 5, which exclusively utilized attention mechanism, eliminating the need for RNNs. They introduced a novel attention mechanism called self-attention, which is also known as KQV-attention (Key, Query, and Value). This attention mechanism subsequently gained prominence as a central component within the Transformer architecture. The attention mechanism stands out due to its ability to provide Transformers with an extensive long-term memory. In the Transformer model, it becomes possible to focus on all previously generated tokens.
Self-attention is a mechanism that allows an input sequence to process itself in a way that each position in the sequence can attend to all positions within the same sequence. This mechanism is a cornerstone of the Transformer architecture, which has revolutionized NLP and beyond by enabling models to efficiently handle sequences of data with complex dependencies. The purpose of self-attention is to compute a representation of each element in a sequence by considering the entire sequence, thereby capturing the contextual relationships between elements regardless of their positional distance from each other. This ability to capture both local and global dependencies makes self-attention particularly powerful for tasks such as machine translation, text summarization, and sequence prediction, where understanding the context and the relationship between words or elements in a sequence is crucial [5]
The mathematical formulation of self-attention involves several key steps. First, a set of query vectors $\boldsymbol{Q}=\boldsymbol{X}\boldsymbol{W}^{Q}$, a set of key vectors $\boldsymbol{K}=\boldsymbol{X}\boldsymbol{W}^{K}$, and a set of value vectors $\boldsymbol{V}=\boldsymbol{X}\boldsymbol{W}^{V}$ are calculated through linear transformations of the input sequence, where $\boldsymbol{X}$ is the input matrix representing embeddings of tokens in a sequence, $\boldsymbol{W}^{Q}$,$\boldsymbol{W}^{K}$, and $\boldsymbol{W}^{V}$ are weight matrices for queries, keys, and values, respectively. The attention scores are then calculated by taking the dot product of the query vector with all key vectors, followed by scaling the result by the inverse square root of the dimension of the keys ($\sqrt{d_{k}}$) to avoid overly large values. These scores are then passed through a softmax function to obtain the attention weights, which represent the importance of each element’s contribution to the output. Finally, the output for each position is computed as a weighted sum of the value vectors,
$\boldsymbol{A(\boldsymbol{Q,K,V})}=\text{softmax}(\frac{\boldsymbol{QK^{\top}}}{\sqrt{d_{k}}})\boldsymbol{V}$ (12)
@@ 2.1.2 Self-Attention
Figure 5: Transformer architecture and its self-attention mechanism, adapted from [5].
@@ 2.1.3 Multi-Head-Attention
Multi-head attention is an extension of the self-attention mechanism designed to allow the model to jointly attend the information from different representation subspaces at different positions [5]. Instead of performing a single attention function, it runs the attention mechanism multiple times in parallel. The outputs of these independent attention computations are then concatenated and linearly transformed into the expected dimension. The mathematical formulation of multi-head attention can be described in four steps. First, Multi-Head Attention Computation,
$\boldsymbol{Q}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{Q},\quad\boldsymbol{K}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{K},\quad\boldsymbol{V}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{V}$ (13)
where i is the index of each self-attention head,
$\boldsymbol{A}_{i}(\boldsymbol{Q}_{i},\boldsymbol{K}_{i},\boldsymbol{V}_{i})=\text{softmax}\left(\frac{\boldsymbol{Q}_{i}\boldsymbol{K}_{i}^{\top}}{\sqrt{d_{k}}}\right)\boldsymbol{V}_{i}$ (14)
Then, concatenating outputs from all heads followed by a final linear transformation,
$\boldsymbol{A}(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\text{Concat}(\boldsymbol{A}_{1},\boldsymbol{A}_{2},\ldots,\boldsymbol{A}_{h})\boldsymbol{W}^{O}$ (15)
where $\boldsymbol{W}^{O}$ is the weight matrix for the final linear transformation.
The multi-head attention mechanism enables the model to capture different types of information from different positions of the input sequence. By processing the sequence through multiple attention ”heads”, the model can focus on different aspects of the sequence, such as syntactic and semantic features, simultaneously. This capability enhances the model’s ability to understand and represent complex data, making multi-head attention a powerful component of Transformer-based architectures [28].
@@ 2.2 From Transformer to Vision Transformer
The journey from the inception of the Transformer model to the development of the Vision Transformer marks a pivotal advancement in deep learning, showcasing the adaptability of models initially designed for sequence data processing to the realm of image analysis. This transition underscores a significant shift in approach, from conventional image processing techniques to more sophisticated, sequence-based methodologies.
Introduced in 2017 by Vaswani et al. [5] through the seminal paper ”Attention Is All You Need,” the Transformer model revolutionized NLP by leveraging self-attention mechanisms. This innovation allowed for the processing of sequences of data without the reliance on recurrent layers, facilitating unprecedented parallelization and significantly reducing training times for large datasets. The Transformer’s success in NLP sparked curiosity about its potential applicability across different types of data, including images, setting the stage for a transformative adaptation.
The adaptation of Transformers for image data pivoted on a novel concept: treating images not as traditional 2D arrays of pixels but as sequences of smaller, discrete image patches. This approach, however, faced computational challenges due to the self-attention mechanism’s quadratic complexity with respect to input length. The breakthrough came with the introduction of the ViT by Dosovitskiy et al. [6] in 2020, which applied the Transformer architecture directly to images, see Figure 6. By dividing an image into fixed-size patches and processing these patches as if they were tokens in a text sequence, ViT was able to capture complex relationships between different parts of an image using the Transformer’s encoder.
@@
For tasks like image classification, the output from the Transformer encoder is passed through a classification head, often utilizing a learnable ”class token” appended to the sequence of patch embeddings for this purpose. The model is trained on large datasets using backpropagation and, during inference, processes images through these steps to predict their classes.
@@
Figure 6: The Vision Transformer architecture, adapted from [6].
2.2.1 Vision Transformer: From Spatial to Spatio-Temporal Tasks
@@
Temporal Dimension Integration The integration of the temporal dimension is a fundamental step in adapting ViT for action recognition. Traditional ViT models process images as a series of patches, treating them essentially as sequences for the self-attention mechanism to analyze spatial relationships. By extending this concept to include the temporal dimension, the models can now treat videos as sequences of frame patches over time. This allows the models to capture the evolution of actions across frames. The work by Bertasius et al. [29] highlights the potential of incorporating temporal information into Transformers, marking a significant advancement in video analysis capabilities.
@@
Spatio-Temporal Embeddings To effectively capture the dynamics of actions within videos, adapted ViT models generate spatio-temporal embeddings. This involves extending the traditional positional embeddings used in ViTs to also include temporal positions, thereby creating embeddings that account for both spatial and temporal information within video sequences. The discussion by Arnab et al. [30] on the creation of these spatio-temporal embeddings showcases the method’s effectiveness in enhancing the model’s understanding of action dynamics across both space and time.
@@
Multi-Head Self-Attention Across Time The extension of self-attention mechanisms to analyze relationships between patches not just within individual frames but also across different frames is crucial for recognizing actions over time. This approach enables the model to identify relevant features and changes across the video sequences, facilitating a deeper understanding of motion and the progression of actions. The exploration by Bertasius et al. [29] of this concept demonstrates how Transformers can be effectively adapted to capture the temporal dynamics of actions, a key aspect of video analysis.
2.3 CNNs: From Spatial to Spatio-Temporal Tasks
CNNs have undergone a significant evolution since their introduction in the 1980s. The Neocognitron [31], developed by Kunihiko Fukushima, presented an early example of neural networks incorporating convolutional operations for image processing, setting the foundations for subsequent progress. Shortly after, Yann LeCun and collaborators introduced LeNet-5 [32], a key architecture designed for handwritten digit recognition, showcasing the effectiveness of convolutional layers in pattern recognition tasks.
The progress of CNNs reached a turning point in the mid-2010s with the introduction of models like AlexNet [33], showcasing their potential in image classification tasks. Alongside architectural innovations, this milestone was achieved thanks to access to large datasets, notably, ImageNet [34], and computational improvements, including the rise of Graphics Processing Units (GPUs) for parallel computing. Large-scale datasets provided the diversity and complexity necessary for training deep networks, while enhanced computational power accelerated the training of sophisticated CNN architectures.
Transfer Learning. Deep Neural Networks, specifically CNNs, have been proven to work very well in the presence of large, diverse datasets. However, access to large representative datasets is a big challenge in the deep learning era. To benefit from the impressive pattern recognition capability of CNNs in the absence of diverse and large datasets, transfer learning comes into play to leverage knowledge gained from a source task to enhance the performance of a model on a related target task. Transfer learning involves pre-training a model on a large dataset through which CNNs’ early layers capture relatively general features and patterns that may be common and helpful for other related target tasks. By freezing the early layers to preserve the learned knowledge, and fine-tuning the later layers on a target task to allow the model to adapt to target task-specific features, the power of deep networks is exploited even when there is limited data. ImageNet is a widely recognised source dataset within the domain of computer vision. Several prominent architectural designs, including VGG, ResNet and GoogleNet have made their pre-trained weights on the Imagenet dataset publicly available to provide a source for subsequent researchers and practitioners. The availability of pre-trained weights accelerates the training process on diverse computer vision tasks, enabling the transfer of learned features and representations from the ImageNet dataset to target applications.
Object detection, another fundamental task in computer vision, has seen significant advancements in the last 10 years. Early methods such as Viola-Jones [38] employed handcrafted features and classifiers to detect and classify objects in a given image. The paradigm shift in the 2010s was marked by the introduction of R-CNN [39]: Region with CNN features and subsequent architectures like Fast R-CNN [39] and Faster R-CNN [40]. This pivotal development brought forth a two-stage approach, involving the proposal of regions of interest and their classification. Following that Single Shot MultiBox Detector (SSD) [41] and You Only Look Once (YOLO) [42] models that allow real-time object detection were proposed to show the efficacy of single-shot detection frameworks. One of the recent aims in the field is to achieve anchor-free methods [43, 44] to work with more flexible and efficient methods.
Segmentation. Similar to object detection, the early techniques for segmentation, a computer vision task that aims to partition an image into distinct and meaningful regions, involved handcrafted features. Early methods like thresholding and region-growing struggled with complex scenes and variations. Advancements in CNNs in the 2010s also skyrocketed the number of proposed architectures that use powerful CNNs and state-of-the-art performance was quickly surpassed by methods like Fully Convolutional Networks (FCN) [45] and U-Net [46]. FCN allowed pixel-level predictions while maintaining spatial information and U-Net helped with better feature integration with the introduction of skip connections. Skip connections, also known as shortcut connections or residual connections, were incorporated in the U-Net architecture to address the challenge of information loss during downsampling and upsampling operations. These connections enable the direct flow of information between corresponding encoder and decoder layers, facilitating the integration of low-level and high-level features. This design choice enhances the model’s ability to capture fine details and spatial information while maintaining pixel-level predictions, making it especially effective for image segmentation tasks.
Transfer learning is not limited to image classification tasks in the computer vision field. Object detection and segmentation tasks, also greatly benefit from the pre-trained weights to enhance model performance and efficiency. In object detection tasks, architectures like R-CNN [39], YOLO [42] and Mask R-CNN [47]; in segmentation tasks architectures like FCN [45] and PSPNet [48] trained on large datasets are employed as a starting point and fine-tuning happens on target tasks. In essence, transfer learning in deep learning, and particularly in CNNs, facilitates the transfer of knowledge, allowing models to achieve impressive results in diverse applications, even when faced with limited labelled data for a target task.
3 Deep Learning-Based Techniques in HAR

### v2

@@ 2.2.4 Types of RNNs
The one-to-one is considered the simplest form of RNNs, where a single input corresponds to a single output. It operates with fixed input and output sizes, functioning similarly to a standard neural network. One-to-many represents a specific category of RNNs that is characterized by its ability to produce multiple outputs based on a single input provided to the model. This type of RNN is particularly useful in applications like image captioning, where a fixed input size results in a series of data outputs. Many-to-one RNNs merge a sequence of inputs into a single output through a series of hidden layers that learn relevant features. An illustrative instance of this RNN type is sentiment analysis, where the model analyzes a sequence of text inputs and produces a single output indicating the sentiment expressed in the text.
@@
Many-to-many RNNs are employed to generate a sequence of output data from a sequence of input units. It can be categorized into two subcategories: equal size and unequal size. In the equal size subcategory, the input and output layers have the same size, see many-to-many architecture in Figure 3c. Several research efforts have emerged to tackle the limitation of the fixed-size input-output sequences in machine translation tasks, as they fail to adequately represent real-world requirements. The unequal size subcategory can handle different sizes of inputs and outputs. A practical application of the unequal size subcategory can be observed in machine translation. In this scenario, the model generates a sequence of translated text outputs based on a sequence of input sentences. Unequal size subcategory employs an encoder-decoder architecture, where the encoder adopts the many-to-one architecture, and the decoder adopts the one-to-many architecture. One notable contribution in this area was made by [51], who pioneered the approach of mapping the entire input sentence to a vector. This work is related to the study conducted by [17], although the latter was specifically utilized to refine hypotheses generated by a phrase-based system [87]. In this architecture, the encoder component plays a crucial role in transforming the inputs into a singular vector, commonly referred to as the context. This context vector, typically with a length of 256, 512 or 1024, encapsulates all the pertinent information detected by the encoder from the input sentence, which serves as the translation target, see Figure 4(a). Subsequently, this vector is passed on to the decoder, which generates the corresponding output sequence. It is important to note that both the encoder and decoder components in this architecture are RNNs. Different from Figure 4(a), Figure 4(b) gives the encoder-decoder architecture with attention which will be introduced in the next section.
@@
Figure 4: Sequence-to-sequence RNN with and without the attention mechanism.
2.2.5 Attention
The evolution of attention mechanisms in neural networks represents a significant advancement in the field of deep learning, particularly in tasks related to NLP and machine translation. Initially introduced by [39], the concept of attention mechanisms was designed to enhance the model’s ability to focus on specific parts of the input sequence when generating an output, mimicking the human ability to concentrate on particular aspects of a task. This foundational work laid the groundwork for subsequent developments in attention mechanisms, providing a mechanism for models to dynamically assign importance to different parts of the input data.
Building on Graves’ initial concept, [6] introduced the additive attention mechanism, which was specifically designed to improve machine translation. This approach computes the attention weights through a feed-forward neural network, allowing the model to consider the entire input sequence and determine the relevance of each part when translating a segment. This additive form of attention significantly improved the performance of sequence-to-sequence models by enabling a more nuanced understanding and alignment between the input and output sequences [87]. Following this, [69] proposed the multiplicative attention mechanism, also known as dot-product attention, which simplifies the computation of attention weights by calculating the dot product between the query and all keys. This method not only streamlined the attention mechanism but also offered improvements in computational efficiency and performance in various NLP tasks, marking a pivotal moment in the evolution of attention mechanisms from their inception to more sophisticated and efficient variants.
@@ 2.2.5 Attention
Figure 5: Transformer architecture and its self-attention mechanism (adapted from [96]).
2.2.6 Self-Attention
To this point, attention mechanisms in sequence-transformation models have primarily relied on complex RNNs, featuring an encoder and a decoder, the most successful models in language translation yet. However, [96] introduced a simple network architecture known as the Transformer, see Figure 5, which exclusively utilized attention mechanism, eliminating the need for RNNs. They introduced a novel attention mechanism called self-attention, which is also known as KQV-attention (Key, Query, and Value). This attention mechanism subsequently gained prominence as a central component within the Transformer architecture. The attention mechanism stands out due to its ability to provide Transformers with an extensive long-term memory. In the Transformer model, it becomes possible to focus on all previously generated tokens.
Self-attention is a mechanism that allows an input sequence to process itself in a way that each position in the sequence can attend to all positions within the same sequence. This mechanism is a cornerstone of the Transformer architecture, which has revolutionized NLP and beyond by enabling models to efficiently handle sequences of data with complex dependencies. The purpose of self-attention is to compute a representation of each element in a sequence by considering the entire sequence, thereby capturing the contextual relationships between elements regardless of their positional distance from each other. This ability to capture both local and global dependencies makes self-attention particularly powerful for tasks such as machine translation, text summarization, and sequence prediction, where understanding the context and the relationship between words or elements in a sequence is crucial [96].
The mathematical formulation of self-attention involves several key steps. First, a set of query vectors $\boldsymbol{Q}=\boldsymbol{X}\boldsymbol{W}^{Q}$, a set of key vectors $\boldsymbol{K}=\boldsymbol{X}\boldsymbol{W}^{K}$, and a set of value vectors $\boldsymbol{V}=\boldsymbol{X}\boldsymbol{W}^{V}$ are calculated through linear transformations of the input sequence, where $\boldsymbol{X}$ is the input matrix representing embeddings of tokens in a sequence, and $\boldsymbol{W}^{Q}$,$\boldsymbol{W}^{K}$, and $\boldsymbol{W}^{V}$ are weight matrices for queries, keys, and values, respectively. The attention scores are then calculated by taking the dot product of the query vector with all key vectors, followed by scaling the result by the inverse square root of the dimension of the keys (say $\sqrt{d_{k}}$) to avoid overly large values. These scores are then passed through a softmax function to obtain the attention weights, which represent the importance of each element’s contribution to the output. Finally, the output say $\boldsymbol{A}$ is computed as a weighted sum of the value vectors, i.e.,
$\boldsymbol{A(\boldsymbol{Q,K,V})}=\text{softmax}(\frac{\boldsymbol{QK^{\top}}}{\sqrt{d_{k}}})\boldsymbol{V}.$ (12)
@@ 2.2.7 Multi-Head-Attention
Multi-head attention is an extension of the self-attention mechanism designed to allow the model to jointly attend the information from different representation subspaces at different positions [96]. Instead of performing a single attention function, it runs the attention mechanism multiple times in parallel. The outputs of these independent attention computations are then concatenated and linearly transformed into the expected dimension. The mathematical formulation of the multi-head attention can be described in the following steps. First, for the $i$-th self-attention head, find
$\boldsymbol{Q}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{Q},\quad\boldsymbol{K}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{K},\quad\boldsymbol{V}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{V},$ (13)
and then compute
$\boldsymbol{A}_{i}(\boldsymbol{Q}_{i},\boldsymbol{K}_{i},\boldsymbol{V}_{i})=\text{softmax}\left(\frac{\boldsymbol{Q}_{i}\boldsymbol{K}_{i}^{\top}}{\sqrt{d_{k}}}\right)\boldsymbol{V}_{i}.$ (14)
The multi-head attention is obtained by concatenating all $\boldsymbol{A}_{i}(\boldsymbol{Q}_{i},\boldsymbol{K}_{i},\boldsymbol{V}_{i})$.
The multi-head attention mechanism enables the model to capture different types of information from different positions of the input sequence. By processing the sequence through multiple attention “heads”, the model can focus on different aspects of the sequence, such as syntactic and semantic features, simultaneously. This capability enhances the model’s ability to understand and represent complex data, making multi-head attention a powerful component of Transformer-based architectures [25].
@@ 2.3 From Transformer to Vision Transformer
The journey from the inception of the Transformer model to the development of the ViT marks a pivotal advancement in deep learning, showcasing the adaptability of models initially designed for sequence data processing to the realm of image analysis. This transition underscores a significant shift in approach, from conventional image processing techniques to more sophisticated sequence-based methodologies.
Introduced by [96] through the seminal paper “Attention Is All You Need”, the Transformer model revolutionized NLP by leveraging self-attention mechanisms. This innovation allowed for the processing of sequences of data without the reliance on recurrent layers, facilitating unprecedented parallelization and significantly reducing training times for large datasets. The Transformer’s success in NLP sparked curiosity about its potential applicability across different types of data, including images, setting the stage for a transformative adaptation.
The adaptation of Transformers for image data pivoted on a novel concept: treating images not as traditional 2D arrays of pixels but as sequences of smaller and discrete image patches. This approach, however, faced computational challenges due to the self-attention mechanism’s quadratic complexity with respect to input length. The breakthrough came with the introduction of the ViT by [29], which applied the Transformer architecture directly to images, see Figure 6. By dividing an image into fixed-size patches and processing these patches as if they were tokens in a text sequence, ViT was able to capture complex relationships between different parts of an image using the Transformer’s encoder.
@@
For tasks like image classification, the output from the Transformer encoder is passed through a classification head, often utilizing a learnable “class token” appended to the sequence of patch embeddings for this purpose. The model is trained on large datasets using backpropagation and, during inference, processes images through these steps to predict their classes.
@@
Figure 6: The ViT architecture (adapted from [29]).
@@
Temporal dimension integration. The integration of the temporal dimension is a fundamental step in adapting ViT for action recognition. Traditional ViT models process images as a series of patches, treating them essentially as sequences for the self-attention mechanism to analyze spatial relationships. By extending this concept to include the temporal dimension, the models can now treat videos as sequences of frame patches over time. This allows the models to capture the evolution of actions across frames. The work by [10] highlights the potential of incorporating temporal information into Transformers, marking a significant advancement in video analysis capabilities.
@@
Spatio-temporal embeddings. To effectively capture the dynamics of actions within videos, adapted ViT models generate spatio-temporal embeddings. This involves extending the traditional positional embeddings used in ViTs to also include temporal positions, thereby creating embeddings that account for both spatial and temporal information within video sequences. The discussion by [4] on the creation of these spatio-temporal embeddings showcases the method’s effectiveness in enhancing the model’s understanding of action dynamics across both space and time.
@@
Multi-head self-attention across time. The extension of self-attention mechanisms to analyze relationships between patches not just within individual frames but also across different frames is crucial for recognizing actions over time. This approach enables the model to identify relevant features and changes across the video sequences, facilitating a deeper understanding of motion and the progression of actions. The exploration by [10] of this concept demonstrates how Transformers can be effectively adapted to capture the temporal dynamics of actions, a key aspect of video analysis.
3 Literature Review